if the box has a surface area of 214cm squared, find the the value of x. side1 is x+1, side 2 is x+2 and side 3 is x
Answer by venugopalramana(3286)

if the box has a surface area of 214cm squared, find the the value of x. side1 is x+1, side 2 is x+2 and side 3 is x
SURFACE AREA =2*(LB+BH+HL)=214
X(X+1)+(X+1)(X+2)+X(X+2)=214/2=107
X^2+X+X^2+2X+X+2+X^2+2X=107
3X^2+6X-105=0
X^2+2X-35=0
X^2+7X-5X-35=0
X(X+7)-5(X+7)=0
(X+7)(X-5)=0
X=5 SINCE NEGATIVE VALUES ARE NOT RELEVANT